Solve the equation y + 2y - V - 2y = 0 using the method of converting to a linear system of first-order ODEs. Show that th
Answer #1

9 (2y - 9 - 2y = 0 Let es 6 , - 9 ใ% - 9 = X2 xon => 2 - 9 Z- 9 29 - 9 - - 27+y + 23 x -22%+ X2+24 • 44 2) - 23 2 4 + 220 - Ri -1 -1 I-3 R=R3-2R 2 | -3 2 Rag = Rz +BR2 Lo o Hence - 1 - 9 = 0 - + W = 0 > u = 9 => W = U 393 315) - v[] ..correspond2 0 2 1 2 1 0 R₂ = R3-R1 2 1 0 0 2 ) 20+1 = 0 => U=- 21+W = 0 =) W--21 | 1-7 2012 ] all 2 - 20 4 corresponding eigenvecter -2